# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 7.3 Dilute nonequilibrium states
The production of pion-enriched matter occurs in ultrarelativistic heavy-ion collisions at highest RHIC and LHC energies. At expansion of the system, the temperature of the pion gas decreases. At a certain temperature, $T_{chem}$, the particle populations stop changing, reaching a state called chemical freeze-out, though the particles still remain in the thermal equilibrium. Then, at a later stage of the expansion up to a thermal freeze-out, till $T(t)>T_{th}$, we are dealing with a system with a dynamically fixed number of pions. At $T<T_{BEC}$, provided $T_{th}<T_{BEC}$, there may appear the Bose-Einstein condensate of pions characterized by the dynamically fixed particle number  Voskresensky (1994); Kolomeitsev et al. (2021).
@@
Another interesting situation is possible: in peripheral heavy-ion collisions, the colliding nuclei can be considered as almost freely interpenetrating beams of nucleons. At the collision energy $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}300$ MeV per nucleon, the nucleon Fermi spheres from each beam do not overlap in momentum space. Since the pion polarization operator is $\Pi\propto-\nu p_{\rm F}k^{2}$, as discussed in Section 5, this corresponds to a density which is effectively four times higher than in the case of overlapping Fermi spheres. Thus, there is a possibility of observing the effects of the $p$ wave pion condensation in peripheral heavy-ion collisions Pirner and Voskresensky (1995); Voskresensky (2025).
7.4 Dilute quantum droplets
@@
The phenomenon of Efimov states in Fermi systems opens another interesting possibility: For particular values of the diluteness parameter $n|a|^{3}$, where $n$ is the fermion density, $a$ is the scattering length, a three-body interaction can provide a dominant contribution in the energy-density functional Bulgac (2002). if this happens, both bosonic and fermionic systems of various sizes and densities could become self-bound. In this relation, we note that the nucleon-nucleon scattering length has an anomalously large magnitude, $a_{NN}\simeq-20$ fm. A discussion of the role of this circumstance for description of a dilute neutron matter can be found in Kolomeitsev et al. (2024). Finally, we notice that some experiments claimed an observation of a tetraneutron ($n^{4}$) signal, which may be associated with a bound state or an unbound resonance state, see discussion in Faestermann and Gernhäuser (2025).
@@
As we have mentioned, in the spectrum of $\pi^{+}$ in neutron matter in the interval of densities $n_{c}^{\pi^{\pm}}>n>n_{c}^{\pi^{+}}$, where $n_{c}^{\pi^{+}}<n_{0}$ according estimate of Migdal et al. (1974), and $n_{c}^{\pi^{\pm}}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}n_{0}$, there appears the branch with $\omega^{\pi^{+}}(k)<0$ at $k\neq 0$. Appearance of such a branch may result in some important consequences, see Voskresensky (1993). For instance, at the formation of the neutron star in a supernova explosion a part of the initial angular momentum can be absorbed by the $\pi^{+}$ condensate subsystem. Metastable $\pi^{+}$-multi-neutron droplets with a large angular momentum and a density $n>n_{c}^{\pi^{+}}$ can be formed. They may take away part of the total angular momentum of the star. Metastable $\pi^{+}$ condensate neutron-rich droplets also can be formed with some probability during heavy-ion collisions. Since density in such droplets might be smaller than $n_{0}$ they may behave as anomalons mentioned above, see Voskresensky (1993). Also there a possibility of appearance of condensate fragments with a large angular momentum in peripheral heavy-ion collisions was discussed.
8 Observations
@@
Search of abnormal nuclei, nuclearites and nuclei-stars composed of absolutely stable matter is an intriguing prospect in astrophysics. Detection of such objects is not a trivial problem due to numerous observational signatures, which are superimposed on those of ordinary neutron stars and white dwarfs. Distinguishing features may be found in ultra-fast rotation (well below 1 ms), star radii less than 9-10 km, specific photon emission from a star surface, etc. Possible candidates among compact objects were identified: RX J1856.5-3754, PSR J1614-2230, PSR J0348+0432, 4U 1820-30, HESS J1731-347, XTE J1814-138. However, definitive conclusions whether these objects are anomalous or not, could not be drawn. Additionally, particular measurements of the low-mass and small-radius compact star HESS J1731-347 depend heavily on the assumed atmosphere model and are sensitive to errors in distance and inferred age. Let us add here the latest results of NICER Mauviard et al. (2025) for PSR J0614-3329 giving the rather small equatorial radius $10.29^{+1.01}_{-0.86}$ km for the mass $M=1.44^{+0.06}_{-0.07}\,M_{\odot}$ (median values with equal-tailed $68\%$ credible interval).
@@
Experimental search for abnormal nuclei was carried out in natural samples among the products of interactions between high-energy particles and matter, in fission products, in heavy-ion collisions, etc., see Refs. Frankel et al. (1976); Holt et al. (1976); Aleshin et al. (1976); Kulikov and Pontecorvo (1977); Price and Stevenson (1975); Frankel and Stevenson (1976); Karnaukhov and Polikarpov (1977); Buchwald et al. (1981). Although no abnormal states were found presently, a number of observations demonstrate anomalies that defy conventional physics interpretations. May be at least some of them are related to the anomalous nuclear states. We mention several such anomalies. A very heavy particle was detected passing through a balloon-borne stack of Cherenkov film, emulsion, and Lexan sheets Price et al. (1975). It could be a neutron anomalous nucleus. The spectra of nuclei obtained in the aerostatic JACEE experiment with an emulsion chamber show a peculiar bend at the cosmic ray energy of 3 PeV. The change in the slope of the spectrum was suggested in Ref. Shaulov et al. (2022) to be explained by the appearance of an anomalous component of cosmic rays, which is much heavier than the nuclear component and, possibly, consists of particles of strange quark matter. In the transition region, both nuclei and strangelets can exist, that leads to large fluctuations in the events observed in the experiments with X-ray emulsion chambers.
@@
Recent works Bertolucci et al. (2017); Zioutas et al. (2020, 2023) argued that there are other observations, which can be hardly explained in the framework of conventional physics. These anomalies include: unexpected correlations with temperature variations in stratosphere, the total electron content of the Earth atmosphere, correlations between earthquake activity and relative positions of planets, unexpected seasonal stratospheric temperature variations, ionospheric anomalies, and their correlations with seismic activity, as well as solar phenomena like the coronal heating problem, the origin of sunspots, and the trigger mechanism of the solar flares, see also Maroudas (2022); Zioutas et al. (2022). The statistical significance of the observed correlations is high. Reference Zhitnitsky and Maroudas (2025) suggested that in mentioned events one may deal with hypothetical very dense and microscopically large composite objects with the mass of the order of grams and sub-micrometer size consisting of nuggets of strongly interacting matter. For instance, it was suggested to look for so-called solar nano-flares with the axion quark nugget annihilation events in the solar corona. Note that, if instead of quark nuggets one considered the Lee-Wick anomalous state or the Migdal pion condensate state abnormal nuclei, at least a part physical consequences would be similar to those for strangelets.
@@
Let us now mention another possibilities. In a uniform electric field with the strength of $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{4}$ V/cm the depth of the potential $V_{0}$ exceeds $2m_{e}\approx 10^{6}$ eV when the distance between plates is $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10$ cm. For a 300  times larger distance between plates $V_{0}$ exceeds two pion masses, $2m_{\pi}\approx 280$ MeV. For comparison, the breakdown voltage for the dry air at atmospheric pressure is $(2\mbox{--}3)\times 10^{4}$ V/cm and for some types of glasses it reaches $3\times 10^{6}$ V/cm. Similar estimates can be found for the cylindrical and spherical capacitors. However we must stress that despite the potential inside the capacitor may readily exceed $2m_{e}$ and $2m_{\pi}$ respectively, the probability of the spontaneous pair production from the vacuum is exponentially small in both cases, $\propto e^{-m^{2}/E}$. Nevertheless, electrons and charged pions can be produced near the walls of the capacitor with a larger probability provided the electric potential well is sufficiently deep. In the broad potential well for $\pi^{-}$ in the region, where $V>m_{\pi}$, the $\pi^{-}$ ground state energy level crosses $\omega=0$ and formation of the static $\pi^{-}$ condensate field becomes energetically favorable. In the region with $V>m_{\pi}$ the static $\pi^{+}$ condensate field can be formed. Note that in the presence of a charged condensate, photons become effectively massive particles via the Anderson-Higgs mechanism. Therefore, the condensate region should look dark. This circumstance can help in identifying the effect.
@@
It is worth noticing that deep potential wells can be formed in thunderstorms on Earth. During tens of minutes before a discharge, the electric voltage may reach values of $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{8}$ V at distance $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(1-10)$ km between clouds, cf. Marshall and Stolzenburg (2001). Hence the typical potential well reaches values $V_{0}\gg m_{e}$ at such distances. An excess of neutrons during thunderstorms has been indeed detected, however it was associated with photonuclear reactions Babich (2019). Strong thunderstorms may also occur in atmospheres of other planets such as Jupiter and Venus.
@@
Finally, let us notice that 10 anomalous bursts correlated with lightnings have been observed in thunderstorms in 5 years Abbasi et al. (2017); Okuda (2019). The estimated energy from individual events within the bursts is five to six orders of magnitude higher than the energy estimated by an event rate. The multiple air showers occurred within 1 ms, defying explanation by conventional high-energy cosmic rays. Similar “exotic events” were recorded by the AUGER collaboration. The ANITA experiment detected two anomalous upward-propagating events with non-inverted polarity. The list of anomalous atmospheric events also includes the Multi-Modal Clustering Events observed by HORIZON 10T.
If a nugget of a new phase of matter passes through the thunderstorm clouds with an ultra-high voltage ($\sim 10^{9}$ V) it will be accelerated and its electron shell will be destroyed in possible collisions with a release of a significant energy and baryon production.
9 Concluding remarks

### v2

@@ VII.3 Dilute nonequilibrium “anomalous” states
The production of pion-enriched matter occurs in ultrarelativistic heavy-ion collisions at highest RHIC and LHC energies. At expansion of the system, the temperature of the pion gas decreases. At a certain temperature, $T_{chem}$, the particle populations stop changing, reaching a state called chemical freeze-out, though the particles still remain in the thermal equilibrium. Then, at a later stage of the expansion up to a thermal freeze-out, till $T(t)>T_{th}$, we are dealing with a system with a dynamically fixed number of pions. At $T<T_{BEC}$, provided $T_{th}<T_{BEC}$, there may appear the Bose-Einstein condensate of pions characterized by the dynamically fixed particle number  [285, 153].
@@
Another interesting situation is possible: in peripheral heavy-ion collisions, the colliding nuclei can be considered as almost freely interpenetrating beams of nucleons. At the collision energy $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}300$ MeV per nucleon, the nucleon Fermi spheres from each beam do not overlap in momentum space. Since at low densities, the main contribution to the pion polarization operator is $\Pi_{\pi NN}\propto-\nu p_{\rm F}(n)k^{2}$, $\nu p_{\rm F}\propto(n\nu^{2})^{1/3}$, see Sect. V. This corresponds to a density, which is effectively four times higher than in the case of overlapping Fermi spheres. Thus, there is a possibility of observing the effects of the p-wave pion condensation in peripheral heavy-ion collisions [225, 290].
VII.4 Dilute quantum droplets
@@
The phenomenon of Efimov states in Fermi systems opens another interesting possibility. For particular values of the diluteness parameter $n|a|^{3}$, where $n$ is the fermion density, $a$ is the scattering length, a three-body interaction can provide a dominant contribution in the energy-density functional [47]. if this happens, both bosonic and fermionic systems of various sizes and densities could become self-bound. In this relation, we note that the nucleon-nucleon scattering length has an anomalously large magnitude, $a_{NN}\simeq-20$ fm. A discussion of the role of this circumstance for description of a dilute neutron matter can be found in [151]. Finally, we notice that some experiments claimed an observation of a tetraneutron ($n^{4}$) signal, which may be associated with a bound state or an unbound resonance state, see discussion in [79].
@@
As we have mentioned, in the spectrum of $\pi^{+}$ in neutron matter in the interval of densities $n_{c}^{\pi^{\pm}}>n>n_{c}^{\pi^{+}}$, where $n_{c}^{\pi^{+}}<n_{0}$ according estimate of Ref. [208], and $n_{c}^{\pi^{\pm}}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}n_{0}$, there appears the branch with $\omega^{\pi^{+}}_{s}(k)<0$ at $k\neq 0$, as we have discussed in Sect. IV.2. Appearance of such a branch may result in some important consequences, see Ref. [284]. For instance, at the formation of the neutron star in a supernova explosion a part of the initial angular momentum can be absorbed by the $\pi^{+}$ condensate subsystem. Metastable $\pi^{+}$-multi-neutron droplets with a large angular momentum and a density $n>n_{c}^{\pi^{+}}$ can be formed, as it was discussed in Sect. III.3. They may take away part of the total angular momentum of the star. Metastable $\pi^{+}$ condensate neutron-rich droplets also can be formed with some probability during heavy-ion collisions. Since density in such droplets might be smaller than $n_{0}$ they may behave as anomalons mentioned above, see Ref. [284]. Also there a possibility of appearance of condensate fragments with a large angular momentum in peripheral heavy-ion collisions was discussed.
VIII Observations
@@
Search of abnormal nuclei, nuclearites and nuclei-stars composed of absolutely stable matter is an intriguing prospect in astrophysics. Detection of such objects is not a trivial problem due to numerous observational signatures, which are superimposed on those of ordinary neutron stars and white dwarfs. Distinguishing features may be found in ultra-fast rotation (well below 1 ms), star radii less than 9-10 km, specific photon emission from a star surface, etc. Possible candidates among compact objects were identified: RX J1856.5-3754, PSR J1614-2230, PSR J0348+0432, 4U 1820-30, HESS J1731-347, XTE J1814-138. However, definitive conclusions whether these objects are anomalous or not, could not be drawn. Additionally, particular measurements of the low-mass and small-radius compact star HESS J1731-347 depend heavily on the assumed atmosphere model and are sensitive to errors in distance and inferred age. Let us add here the latest results of NICER [200] for PSR J0614-3329 giving rather small equatorial radius $10.29^{+1.01}_{-0.86}$ km for the mass $M=1.44^{+0.06}_{-0.07}\,M_{\odot}$ (median values with equal-tailed $68\%$ credible interval).
Experimental search for abnormal nuclei was carried out in natural samples among the products of interactions between high-energy particles and matter, in fission products, in heavy-ion collisions, etc., see Refs. [89, 124, 7, 167, 232, 88, 141, 14, 13, 13, 15, 46]. Reference [134] accounted for the absence of traces found in the sample previously considered in Ref. [230] and excluded strangelets with a mass $M<55$ g. More details can be found in [67].
@@
Although no abnormal states were found, presently a number of observations demonstrate anomalies that defy conventional physics interpretations. May be at least some of them are related to the anomalous nuclear states. We mention several such anomalies. A very heavy particle was detected passing through a balloon-borne stack of Cherenkov film, emulsion, and Lexan sheets [231]. It could be a neutron anomalous nucleus. The spectra of nuclei obtained in the aerostatic JACEE experiment with an emulsion chamber show a peculiar bend at the cosmic ray energy of 3 PeV. The change in the slope of the spectrum was suggested in Ref. [261] to be explained by the appearance of an anomalous component of cosmic rays, which is much heavier than the ordinary nuclear component and could be associated, e.g., with quark nuggets or abnormal $\pi$ or $\sigma$ condensate nuclei. In the transition region, both nuclei and strangelets can exist, that leads to large fluctuations in the events observed in the experiments with X-ray emulsion chambers.
@@
Recent works [26, 304, 305] argued that there are other observations, which can be hardly explained in the framework of conventional physics. These anomalies include: unexpected correlations with temperature variations in stratosphere, the total electron content of the Earth atmosphere, correlations between earthquake activity and relative positions of planets, unexpected seasonal stratospheric temperature variations, ionospheric anomalies, and their correlations with seismic activity, as well as solar phenomena like the coronal heating problem, the origin of sunspots, and the trigger mechanism of the solar flares, see also [192, 306]. The statistical significance of the observed correlations is high. Reference [301] suggested that in mentioned events one may deal with hypothetical very dense and microscopically large composite objects with the mass of the order of grams and sub-micrometer size consisting of nuggets of strongly interacting matter. For instance, it was suggested to look for so-called solar nano-flares with the axion quark nugget annihilation events in the solar corona. Note that, if instead of quark nuggets one considered the Lee-Wick abnormal states or the Migdal pion condensate abnormal nuclei, at least a part of physical consequences would be similar to those for strangelets.
@@
We now mention another possibilities. In a uniform electric field with the strength of $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(10^{4}-10^{5})$ V/cm the depth of the potential $V_{0}$ exceeds $2m_{e}\approx 10^{6}$ eV when the distance between plates is $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10$ cm. For a 300  times larger distance between plates $V_{0}$ exceeds two pion masses, $2m_{\pi}\approx 280$ MeV. For comparison, the breakdown voltage for the dry air at atmospheric pressure is $(2\mbox{--}3)\times 10^{4}$ V/cm and for some types of glasses it reaches $3\times 10^{6}$ V/cm. Similar estimates can be found for the cylindrical and spherical capacitors. However we must stress that despite the potential inside the capacitor may readily exceed $2m_{e}$ and $2m_{\pi}$ respectively, the probability of the spontaneous pair production from the vacuum is exponentially small in both cases, $\propto e^{-m^{2}/E}$. Nevertheless, electrons and charged pions and, maybe, kaons can be produced with the larger probability near the walls of the empty capacitor or in reactions, if capacitor is filled by a medium, provided the electric potential well is sufficiently deep. In the broad potential well for $\pi^{-}$ in the region, where $V>m_{\pi}$, the $\pi^{-}$ ground state energy level crosses $\omega=0$ and formation of the static $\pi^{-}$ condensate field becomes energetically favorable. In the region with $V>m_{\pi}$ the static $\pi^{+}$ condensate field can be formed, if the time needed for the formation of the condensate is sufficient. Note that in the presence of a charged condensate, photons become effectively massive particles via the Anderson-Higgs mechanism. Therefore, in the case of empty capacitor the condensate region should look dark. This circumstance can help in identifying the effect.
@@
It is worth noticing that deep potential wells can be formed in thunderstorms on Earth. During tens of minutes before a discharge, the electric voltage may reach values of $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{8}$ V at distance $\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(1-10)$ km between clouds, cf. [193]. Hence the typical potential well reaches values $V_{0}\gg m_{e}$ at such distances. An excess of neutrons during thunderstorms has been indeed detected, however it was associated with photonuclear reactions [18]. Note that pions, which could form a condensate, can also be produced in photonuclear reactions. Strong thunderstorms may also occur in atmospheres of other planets such as Jupiter and Venus. If a nugget of a new phase of the matter passes through the thunderstorm clouds at the ultra-high voltage ($\sim 10^{9}$ V) between them, the nugget will be accelerated and its electron shell will be destroyed with a release of the sufficient energy for the production of baryons, pions, quarks, etc. In these processes the nugget can be destroyed that could be observed via the significant energy release.
@@
We should notice that 10 anomalous bursts correlated with lightnings have been observed in thunderstorms occurred during 5 years [1, 220]. The estimated energy from individual events within the bursts is five to six orders of magnitude higher than the energy estimated by an event rate. The multiple air showers occurred within 1 ms, defying explanation by conventional high-energy cosmic rays. Similar “exotic events” were recorded by the AUGER collaboration. The ANITA experiment detected two anomalous upward-propagating events with non-inverted polarity. The list of anomalous atmospheric events also includes the Multi-Modal Clustering Events observed by HORIZON 10T. Among various possible explanations we indicate the possibility of the energy release due to the destruction of the previously bounded nuggets arrived from cosmos.
IX Conclusion
In 1971 Migdal suggested a possibility of existence of metastable or stable abnormal pion condensate nuclei and Bodmer assumed possibility of collapsed quark nuclei. In 1974 Lee and Wick proposed scalar (sigma) condensate abnormal nuclei. In 1984 Witten suggested quark nuggets. These pioneering suggestions paved the way for numerous subsequent theoretical and experimental investigations of various anomalous nuclear objects. New more detailed models were constructed. The review is dedicated 115-year anniversary of A. B. Migdal and 100-year anniversary of T. D. Lee. Our aim was to present the development of their ideas and discuss the modern status of the subject, demonstrating the interrelation between the different phenomena under discussion.
In Sect. II we reviewed the original Lee-Wick model of scalar condensate abnormal nuclei. The model describes the coupling of nucleons with a scalar mean field. Typical behaviors of the effective potential and the energy of the model as function of a scalar field $\phi$ are shown in Fig. 1 and Fig. 2. The role of the scalar field can be played by the $\sigma$ meson and the model then can be treated as a $\sigma$ model permitting chiral symmetry breaking in a dense nuclear matter. The same year when Lee and Wick proposed their model, Walecka formulated a relativistic mean-field $\sigma-\omega$ model [295, 52] with two coupling constants fitted to describe the nuclear binding energy and the value of the nuclear saturation density. Various modifications of the Walecka model were introduced for the better description of nuclear matter properties. The density dependencies of the effective nucleon mass within the Lee-Wick model and in the RMF model are shown in Fig. 3. In the Lee-Wick model the effective nucleon mass undergoes a jump to a tiny value at the critical density, whereas it decreases smoothly with the density increase in the standard RMF models. With parameters, which were fitted to describe the nuclear system at $n=n_{0}$, at larger densities in these RMF models the Lee-Wick abnormal state does not appear. However information about chiral symmetry is lost in the standard RMF based models, whereas it is present in the Lee and Wick model. Although not permitting to describe properties of the ordinary atomic nuclei and nuclear matter at $n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n_{0}$, the Lee-Wick model demonstrates a principle possibility of existence of the new stable (or metastable) abnormal nuclear state at $n$ significantly above $n_{0}$. As a possibility to reconcile these approaches, we indicated the RMF models employing the $\sigma$ field scaled hadron masses and coupling constants [154, 143, 144, 195, 196], which can be treated as models using the idea of the partial chiral symmetry conservation. For simplicity the parameters of the models used in [154, 143, 144, 195, 196] were chosen such that the new Lee-Wick minimum is not realized. Changing parameters of the models we could get new solutions, some of them with a jump to a low value of the effective nucleon mass at a critical density. Such a possibility is realized in the model [152] at some choice of the attractive $\Delta$ optical potential, see Fig. 9 in Sect. V. This question should be more carefully studied in the future.
Section III considered various aspects of the model of supercharged nuclei, nuclearites and nuclei-stars glued together by light charged bosons, if the latter existed. In reality the role of the light charged boson can be played by the ordinary charged pion since in the nuclear medium and in external fields the pion effective mass can be essentially decreased. For example, it was demonstrated that a rapid rotation of a nugget acts in favor of the decrease of the effective pion mass and appearance of the stable state.
Section IV analysed the behavior of the pion degree of freedom in dense nuclear matter and the possibility of the Migdal’s p-wave pion condensation, cf. Ref. [210]. Equation (25) is the final result of the resummation of the nucleon-nucleon interaction amplitude in the spin-isospin channel in the dense nuclear matter done within the Fermi liquid approach of Migdal allowing for the explicit treatment of soft (quasi-Goldstone) pion degree of freedom. For $n>n_{0}$ the in-medium pion exchange term in (25) becomes dominant and the nucleon-nucleon scattering cross section appreciably increases with increasing density towards the pion condensation point, see Eq. (31). The $\Delta$ isobar plays very important role because the mass of the $\Delta$ isobar is only $2m_{\pi}$ larger than the nucleon mass and degeneracy factor is 4 times larger. Moreover the pion-Delta-nucleon attraction coupling constant is twice larger than that of the pion-nucleon-nucleon one.
Various types of the p-wave pion condensation ($\pi^{+}$, $\pi^{\pm}$, $\pi^{0}$) as well as a possibility of the s wave pion condensation were mentioned. The typical effective pion mass (better saying the effective pion gap) in the dense medium depends on the pion momentum and decreases with increasing baryon density. It is shown in Fig. 6. The description of the pion degree of freedom can be formulated within the linear $\sigma$-model [49, 50, 22, 280, 210, 276], in which both the pion and sigma field condensations can be treated simultaneously. Also the $\sigma(n)$ dependence is related to the change of the quark condensate in baryon matter. The possibility of self-bound pion condensate abnormal nuclei, nuclearites and nuclei-stars found within the chiral $\sigma$-model of pion condensation used in Ref. [280] is demonstrated in Fig. 8.
Section V reviewed the idea of $\Delta$ resonance matter. Occupation of the new $\Delta$ Fermi sea with increasing baryon density occurs within many models for the equation of state of neutron star matter already at $n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(2-3)n_{0}$. Also, $\Delta$ isobars may appear in symmetric nuclear matter. Appearance of $\Delta$ isobars results in an energy decrease making the nuclear matter more stable. It may result in a metastable or even stable ground state of the $\Delta$ resonance matter if the attractive $\Delta$ optical potential is sufficiently strong. Fig. 9 (left) demonstrates an abrupt decrease of the nucleon effective mass to a small value occurring in the isospin-symmetric matter at a critical density $n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(3-4)n_{0}$ in a RMF model [196] employing the $\sigma$ scaled hadron masses. We note that such a behavior is also typical for the Lee-Wick model.
Section VI reviewed ideas of Bodmer, Witten and others on quark strangelets or nuclearites and strange stars as stable baryonic states composed of deconfined quarks. These states are interconnected with nontopological solitons containing baryon constituents studied by Lee and collaborators as well as by Coleman [90, 55, 184, 178, 177]. The idea of stable strange quark matter by Witten 1984 [298] is somewhat similar to the idea of stable $\Delta$ resonance matter by Troitsky and Chekunaev 1979 [269]. If a model predicts the strange quark matter to be more stable than an iron nucleus, one speaks about strangelets and quark nuclearites, provided the baryon number $A\ll 10^{56}-10^{57}$, and about strange stars for $A\sim 10^{56}-10^{57}$. In the latter case the gravity essentially contributes to the binding. The surface layer of these objects has a typical length of $\sim 10$ fm. Some other models do not allow for the absolutely stable strange quark matter but allow for metastable matter at switching off the gravity. Other models use parameters that render strange and not strange ($u,d$ quark) deconfined matter unstable. Nevertheless, we may deal with the quark matter in interiors of compact stars named hybrid stars. All hybrid stars have macroscopic hadron shells.
Various models describing quark deconfined state were reviewed. Among different models we indicated the heavy quark quasiparticle model [129] capable to describe the lattice data [86, 58]. At some choice of the parameters the model [129] allows for the stable quark matter but does not permit the hybrid stars.
At $T<T_{\Delta}$ for the quark-quark pairing the quark matter should form a color superconductor. We briefly discussed this phenomenon. We focused on a possible manifestation of the pion and $\sigma$ condensate nuclei-stars and strange stars via the measurement of the MRM-R relation and via the cooling of these stars.
Section VII discussed the possibility of the existence of very dilute stable or metastable nuclear systems. For this the Fermi liquid Green’s function approach was employed as for the description of dense nuclear systems. However, in the case of pion-condensate nuclear objects we used nucleon-nucleon interactions in the spin channel (or in spin-isospin one), whereas in Sect. VII we focus on the scalar channel of an $NN$ interaction. In this channel, the interaction is expressed via the scalar Landau-Migdal parameter $f_{0}(n)$. First we showed that in case of approximately isospin-symmetric matter at low densities we have $f_{0}(n)<-1$, and there appears the Pomeranchuk instability. Then we argued in [158] that this instability may result in the formation of a static scalar-field condensate (compare with the Lee-Wick scalar field condensation). We demonstrated that a decrease in the energy due to the condensation can be sufficient to allow for the existence of very dilute metastable or even stable abnormal dilute nuclei held together by the condensate of a scalar field, see Fig. 11.
Then we studied if it is possible to form a Bose-Einstein condensate from bosonic clusters consisting of an even number of nucleons, e.g. of the $\alpha$ clusters or deutrons. We showed that the charged Bose cluster condensates can be described by a complex field $\phi$ similarly to the description of supercharged nuclei performed in Sect. III. We estimated the energy for the dilute $\alpha$ condensate extended systems, which charge is compensated by the electrons. For $n>n_{\rm Mott}\simeq 0.04n_{0}$ according to our estimate the system is destroyed.
Finally we made some speculations about formation of metastable pion Bose-Einstein condensate during the course of the heavy-ion collisions. Section VIII discussed the situation with observations of exotic phenomena considered in the review.